solve for x: sin -1 (1-x) - 2sin -1 x = pi/2

solve for x:


sin-1(1-x) - 2sin-1x = pi/2

Grade:12

2 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear

sin-1(1-x) - 2sin-1x = pi/2

sin-1(1-x)  = pi/2 + 2sin-1x

let x = sinΘ

so

sin-1(1-x)  = pi/2 + 2sin-1sinΘ

1-x = sin(pi/2 + 2Θ)

1-x = cos2Θ

1-x = 1-2sin2Θ

 1-x = 1-2x2

 x=2x2

x(2x-1) =0

x=0 ,1/2

x=1/2 is not satisfying the given equation so

final solution is x=0

Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand - post it here and we will get you
the answer and detailed  solution very  quickly.

 We are all IITians and here to help you in your IIT JEE preparation.

All the best.
 
Regards,
Askiitians Experts
Badiuddin

Kushagra Madhukar
3 years ago

Dear student,
Please find the attached solution to your problem below. sin-1(1-x) - 2sin-1x = pi/2

sin-1(1-x) = pi/2 + 2sin-1x
let, x = sinΘ
so sin-1(1-x) = pi/2 + 2sin-1sinΘ
1-x = sin(pi/2 + 2Θ)
1-x = cos2Θ
1-x = 1-2sin2Θ
1-x = 1-2x2
x=2x2
x(2x-1) =0
x=0 ,1/2
x=1/2 is not satisfying the given equation so
final solution is x=0

Hope it helps.
Thanks and regards,
Kushagra